OpenStudy (anonymous):

7x - 2(x + 1) = 6x + 14 plzz Help !

OpenStudy (anonymous):

7x - 2(x + 1) = 6x + 14 7x - 2x -2 = 6x +14 7x - 6x -2x -2 -14 = 0 -x -16 = 0 -x = 16 x = -16

OpenStudy (anonymous):

just easy

OpenStudy (anonymous):

Break it down So your equation 7x - 2 (x+1) = 6x +14 Step 1 : 7x-2(x+1) = 6x+14 Step 2: 7x-2x-2 = 6x+14 Expand your equation Step 3: 5x-2 = 6x+14 combine your like terms Step 4: -2 = x+14 subtract 5x Your answer will look like this : Then subtract 14 x = -16 Hope this helped!